• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name zhangyi

    Log entry time 08:03:26 on March 9,2009

    Entry number 265417

    keyword=owl shift summary


    Owl shift summary: Yi Zhang (SL), Wolfgang Troth (TO), Miha Mihovilovic (3rd)
    =========================================================================

    EVENTS
    ------

    02:14 restarted both Happex and compton run, finished a NMR measurement
    before start new runs in Bigbite and LHRS
    02:45 MCC told us we would lost beam for vaccue problem. After 10 min waiting
    we stoped both runs.
    03:16 MCC called again and said that the beam would lost for at least 3 hours.


    RUNS
    ----

    Bigbite:

    23:58-01:00 run #2083 prod 5M
    01:01-02:03 run #2084 prod 5M
    02:03-02:45 run #2085 prod 1.5M

    HRSL:


    23:58-02:03 #20607 653k prod. stoped for NMR measurement
    02:03-02:45 #20608 143k prod. stoped for beam trip